### v1

@@ Abstract
Motivation. Alongside optimized strains and processes, texture is a key quality attribute that shapes consumer acceptance. It arises from a complex interplay of physical, mechanical, and sensory properties and spans multiple scales–from molecular structure to oral processing–making it difficult to quantify and predict. Current methods rely heavily on constly sensory panels and labor-intensive mechanical tests such as cutting tests, tension, compression, and shear tests, food rheology, and texture profile analysis. The challenge is to bridge intuitive descriptors like crispiness, chewiness, or fibrousness with measurable and predictable parameters that ultimately drive product design. Can AI integrate multimodal datasets—from force–deformation curves and rheological measurements—to predict and ultimately engineer food textures that accelerate the acceptance of sustainable proteins?
Figure 6: AI for texture. Machine learning can integrate multimodal data, from ingredients to process conditions, to optimize plant-based meat analogues, from microstructural appearance to macrostructural texture. Future machine learning tools could help us reverse engineer ingredients and process conditions to achieve desired appearance and texture.
@@
Background and State of the Art. Every ingredient and product possesses a characteristic texture, which we can engineer through processing techniques such as extrusion, shear cell structuring, or 3D printing to mimic animal-based foods. Recent advances show the promise of machine learning for decoding texture–structure relationships: constitutive neural networks can map mechanical signatures of plant-based meat alternatives onto interpretable and generalizable physics-based models [68, 69]. Autoencoders can predict sensory texture attributes from rheological data and advance data analysis beyond traditional correlation methods [39]. Classifiers such as support vector machines and neural networks can already identify snack freshness from mechanical and acoustic signals with up to 92% accuracy [70]. Researchers are using machine learning to predict viscosity, gelation, and foaming of plant proteins [71, 34]. Meanwhile, texture analysis from images is rapidly growing. Transfer learning with models like VGGNet, ResNet, and DenseNet allows us to predict the cooking times for fish with potential extension to plant-based foods [72], while ML-driven microscopic image analysis can help us identify droplets in colloidal systems to link microstructure to texture [73]. Beyond academic applications, commercial platforms already harness AI for texture optimization: For example, NotCo [42, 43, 44, 45] and Climax Foods successfully deploy AI to mine ingredient databases, combine plant ingredients, and design formulations that replicate animal textures.
@@
Applications and Opportunities. AI accelerates prototyping by predicting textural outcomes before physical trials and reduces time- and resource-intense trial-and-error cycles. In addition, ingredient design and process optimization can balance target textures with environmental impact, for example, to reduce greenhouse gas emissions. Another important application is personalization, for example designing softer foods for elderly populations to improve swallowing safety. LLMs can also convert consumer feedback such as “too chewy” or “not crispy enough” into quantifiable parameters for product adjustment and bridge the gap between perception and engineering [74]. To address data sparsity, LLMs can generate synthetic training data by pairing realistic texture descriptors with mechanical properties. Ideally, AI can integrate data across multiple length scales, from molecular interactions to macroscopic structure, to capture the hierarchical origins of texture. Figure 6 illustrates how machine learning can integrate multimodal data, from ingredients to process conditions, to predict outcomes across the scales, from microstructural appearance to macrostructural texture. Inversely, future machine learning tools could help us reverse engineer ingredients and process conditions to achieve a desired texture. More broadly, generative design tools, similar to those used for inorganic materials [7], hold the potential to propose ingredient–process combinations that meet desired textural constraints.
@@
Challenges and Open Questions. Significant obstacles remain before AI-driven texture engineering will truly become routine. The largest bottleneck is the lack of large, open datasets that couple textural, sensory, and processing data. Industrial data sharing also remains limited. Data fusion remains technically challenging, as mechanical, rheological, visual, and compositional data are inherently multimodal and often unstructured. Foundation models trained on multiple modalities may help, but progress requires standardized protocols for measurement, evaluation, and reporting. To succeed, the field must establish robust multimodal datasets, shared standards and interoperable platforms. The ultimate vision is an AI system for inverse design–with a specific target texture as input and optimal ingredients and processes as output–that transforms texture development into a predictive, model-guided science.
AI for Sensory Properties
@@
Motivation. Ultimately, food is not only a material design challenge, but a sensory experience. Aroma, flavor, and texture are the strongest determinants of consumer acceptance. The central challenge is to map measurable physical and chemical properties of ingredients to the multidimensional space of human perception. Machine learning offers a powerful framework to learn this mapping, enabling both the prediction of sensory profiles and the computational design of novel ingredients with desired characteristics. How far are AI models from accurately predicting human sensory perception, and from enabling the design of novel flavor and aroma experiences?
Figure 7: AI for sensory properties. Optimizing the sensory perception of a food product is traditionally a costly, time-consuming process involving expert tasting panels and instrumental analysis including mass spectrometry and gas chromatography. In contrast, an AI-powered sensory pipeline has the potential to predict people’s sensory perceptions from only the chemical or molecular make-up of the food, eliminating the need for costly human panels and vastly reducing time spent iterating product formulations.
@@
Background and State of the Art. Traditional sensory science has relied on human panels for perceptual evaluation and on analytical instruments such as gas chromatography–mass spectrometry for chemical composition. Panels remain the gold standard but are slow and expensive, while instruments offer precision without perceptual insight. Early machine learning approaches combined biomimetic sensors, such as E-noses, with classifiers [75], but required heavy feature engineering and were limited to narrow classification tasks [76]. The field has since shifted toward deep representation learning [77], which learns general-purpose sensory maps directly from raw data. The Principal Odor Map exemplifies this trend: a graph neural network organizes molecules into a vector space where distance encodes perceptual similarity, achieving human-level accuracy in odor labeling [22]. Researchers employ attention-based architectures to extend this mapping to molecular mixtures [78, 79]. Figure 7 highlights these advances: Algorithmic tools for predictive sensory modeling are maturing rapidly, but progress is constrained by the scarcity and cost of high-quality perceptual datasets.
@@
Applications and Opportunities. Learned sensory representations open powerful opportunities for food innovation. High-throughput virtual screening of molecular libraries can identify novel flavor compounds for plant-based meats, accelerating product development [11]. Integrating predictive models with low-cost, miniaturized sensors enables real-time quality monitoring in manufacturing and packaging [80], for example with wireless freshness indicators that reduce food waste [81]. These tools also provide systematic frameworks for biomimicry, allowing formulators to predict how ingredient or process modifications alter the final sensory profile. Still, a feedback loop remains essential: computational predictions must be continuously validated and refined with real-world experiments to ensure consumer relevance.
@@
Challenges and Open Questions. Today, the bottleneck is no longer algorithmic capacity but high-quality data acquisition [82, 83]. Deep learning models require large multimodal datasets, yet sensory panels cannot generate data at the necessary scale to capture the vast chemical space of food. The path forward lies in closing the loop between modeling and experiment. Generative models such as transformers and diffusion architectures are algorithmically ready, but their success depends on new data pipelines. Developing robust, low-cost hardware and embedding it into automated platforms will be critical to unlock the full potential of AI for sensory science. The ultimate vision is a Self-Driving Lab that unites synthesis, characterization, and machine learning into platforms to autonomously generate and analyze multimodal data and accelerate de novo food design.
AI for Manufacturing
@@
Motivation. Even the most well-tuned designs will fail if they cannot be produced consistently and affordably. Manufacturing convincing meat analogues is inherently complex because technical goals such as texture, flavor, appearance, and shelf life interact non-linearly with environmental and economic constraints. Relevant datasets span ingredient composition and structure, rheology, thermal behavior, processing steps such as extrusion or fermentation, as well as telemetry, sensory, and consumer data. Through digital twins, computer vision, and other modes, AI is emerging as a key technology to integrate these multimodal datasets and streamline process optimization. Typical model inputs include formulation and process profiles, while outputs can range from texture, color, and flavor to cost and environmental impact. How can we ensure that AI-driven manufacturing pipelines deliver reliable, scalable, and cost-effective production of alternative proteins without compromising sensory quality or sustainability goals?
Background and State of the Art. AI and ML are deployed across the manufacturing pipeline–from formulation and process design to in-line quality control. For example, high-moisture extrusion can generate fibrous, anisotropic structures that mimic muscle tissue, but outcomes are highly sensitive to process parameters such as moisture, temperature, and screw configuration. Bayesian optimization and other data-driven strategies explore this parameter space more efficiently than traditional trial-and-error [85, 86]. Researchers can use informatics to link protein structure and functionality, for example, gelation, emulsification, water or fat binding, to downstream performance. This ultimately allows designers to move beyond commodity proteins, but economic viability remains a major hurdle. The current price gap may be addressed with higher throughput, improved yields, cheaper inputs, and greater capacity utilization [87]. Meanwhile, digital twins show promise for soft sensing, scenario testing, and closed-loop control, though most food-sector deployments remain at pilot scale [88]. Data fragmentation persists as a key obstacle: measurements are dispersed across proprietary plant logs, supplementary materials, and discipline-specific repositories with inconsistent metadata. FAIR data principles tailored to agri-food [89], coupled with large language models for literature-scale information extraction [90], offer pathways to build structured knowledge graphs that connect formulation–process–structure–sensory chains for meat analogs.
@@
Applications and Opportunities. AI is opening multiple opportunities for advancing alternative protein manufacturing. Engineers can use machine learning and Bayesian optimization to tune high-moisture extrusion parameters to target fibrousness, yield, and energy efficiency, while digital twins provide what-if simulations for scale-up and enable real-time sensing [84, 85, 87]. Food scientists can use informatics pipelines to rank plant proteins and design blends at scale using rheology–functionality mappings and literature-derived property tables, allowing ingredient discovery to expand well beyond commodity sources [86, 89]. Large language models further act as formulation copilots, transform unstructured texts into structured knowledge graphs, and suggest experiments to resolve uncertainty [89]. At the same time, computer vision and hyperspectral analytics can advance in-line quality assurance, linking surface color and structure to sensory proxies to enable continuous assessment during production [87]. Together, these approaches illustrate how AI can accelerate prototyping, reduce waste, and bridge the gap between laboratory development and industrial manufacturing.
@@
Challenges and Open Questions. Scaling these tools requires high-quality, shareable datasets that link formulations, processes, structures, and sensory outcomes with standardized metadata. Proprietary and fragmented data currently hinder model transfer and benchmarking, underscoring the need for domain-specific FAIR implementations, shared testbeds, and community baselines [88]. Ingredient variability remains a crucial bottleneck– robust hybrid models are needed to achieve reliable process control [84, 85]. Industrial adoption also demands capital investment, workforce training, and validation on real production lines–most food digital twins have not been tested at scale [87]. The ultimate vision is to establish data standards and LLM-enabled knowledge base that make AI-assisted formulation and control routine in manufacturing for plants to deliver consistent sensory performace at—or below— price parity with conventional meat, with lower cost, waste, and energy demand.
AI for Recipes
@@
Motivation. Beyond production, food preparation determines consumer experience of alternative proteins. When considering this, both individuals and foodservice operations must balance enjoyment with health, cost, convenience, and allergens, while advancing broader goals like protein diversification and sustainability. Large Language Models are promising aids, as they are already trained on vast collections of recipes and can integrate data from sources such as Recipes1M, USDA FoodData Central, the HESTIA climate database, and Food.com reviews. How can we best integrate human knowledge into AI systems to ensure that recipe models generalize across diverse groups of users, dietary needs, and cultural traditions?
Background and State of the Art. Recent research has evaluated and refined LLMs for recipe-related tasks, including generation, revision, and preference prediction. Improvements rely on several strategies: fine-tuning with domain-specific datasets, retrieval-augmented generation that combines LLMs with external knowledge at inference time, and integration with optimization algorithms or external tools. A study benchmarking models on the Recipe1M dataset found off-the-shelf LLMs unreliable due to hallucinations [91], echoing anecdotal reports [92, 93]. However, fine-tuning on Recipe1M produced state-of-the-art results [91, 94], and supervised fine-tuning with direct preference optimization has enabled recipe revision tasks [95]. Other work demonstrates the benefits of retrieval-augmented generation for recipes [96, 97], while one study directly targeted sustainable menu design [98]. Encouragingly, some LLMs now match expert food scientists in preference modeling for alternative protein applications [98]. Combining these models with combinatorial optimization algorithms has the potential to generate recipes and menus that satisfy nutritional and sustainability goals while complementing expert chefs.
@@
Applications and Opportunities. The ability to generate recipes under explicit constraints creates opportunities across individual and institutional contexts. For individuals, fine-tuned LLMs could power AI health coaches that teach users how to modify meals to be more sustainable, nutritious, or culturally appropriate, while respecting restrictions such as allergies or religious requirements. Chefs and students could practice recipe generation in educational applications, receiving AI-driven feedback to refine their skills [99]. In institutional settings, LLMs could support foodservice operations in offices, universities, or hospitals, tailoring descriptions, ingredients, and preparation methods under strict constraints. This is particularly valuable in programs such as “Food as Medicine,” which deliver cost- and time-sensitive meals for patients with chronic conditions after hospital discharge [100]. In all these scenarios, AI could accelerate constrained recipe generation, broaden access to healthy and sustainable foods, and scale public health interventions that rely on tailored meal planning.
@@
Challenges and Open Questions. Despite this promise, important challenges remain. Off-the-shelf LLMs perform unevenly across cultural contexts, often fail to capture diverse user preferences, and can generate outputs that are numerically inconsistent, unsafe, or infeasible. They also lack molecular-level understanding of food and have limited knowledge of food safety. Training LLMs from scratch remains prohibitively resource-intensive [101]. These shortcomings raise risks of culturally inappropriate, impractical, or even dangerous recipes, while incurring environmental costs. Mitigation strategies include equipping LLMs with external datasets, relying on open-source pre-trained models, and carefully integrating optimization tools. Beyond technical concerns, broader societal impacts must be addressed: LLMs should promote user agency and creativity rather than replace them, and they should be designed to complement, not displace, human expertise. The vision is a human-centered approach in which LLMs assist with recipe drafts and revisions while empowering users to adapt meals to their own needs, preferences, and cultures, ultimately supporting sustainable and healthy food choices.
Summary and Outlook

### v2

@@ Results
Texture: Learning structure–perception maps. Texture is a key quality attribute that shapes consumer acceptance. It emerges from coupled physical, mechanical, and sensory properties across scales—from molecular structure to oral processing—which complicates quantification and prediction. We distinguish between physical texture, measured through mechanical and rheological tests, and sensory texture, perceived through attributes such as chewiness or fibrousness. This distinction links measurable material behavior to human perception and defines how AI can connect mechanics with sensory response. Current approaches rely on costly sensory panels and labor-intensive tests such as cutting, tension, compression, shear, and rheology. The core challenge is to map intuitive descriptors like crispiness or chewiness onto measurable parameters that guide product design (Fig. 3B). This raises the question of how AI can integrate multimodal datasets to predict and engineer consumer-preferred textures and accelerate the adoption of sustainable proteins.
@@
Every ingredient and product possesses a characteristic texture, which we can engineer through processing techniques such as extrusion, shear cell structuring, or 3D printing to mimic animal-based foods. Recent advances show the promise of machine learning for decoding texture–structure relationships: Constitutive neural networks can map mechanical signatures of plant-based meat alternatives onto interpretable and generalizable physics-based models [63] that achieve ${\textsf{R}}^{2}$ values from 0.92 to 0.99 for commercially available plant-based meats [64]. Autoencoders can predict sensory texture attributes from rheological data and advance data analysis beyond traditional correlation methods [35]. Classifiers such as support vector machines and neural networks can identify snack freshness from mechanical and acoustic signals with up to 92% accuracy [65]. Researchers are using machine learning to predict viscosity, gelation, and foaming of plant proteins [66, 33]. Meanwhile, texture analysis from images is rapidly growing. Transfer learning with models like VGGNet, ResNet, and DenseNet allows us to predict the cooking times for fish with potential extension to plant-based foods [67], while ML-driven microscopic image analysis can help us identify droplets in colloidal systems to link microstructure to texture [68]. Beyond academic applications, commercial platforms already harness AI for texture optimization: For example, NotCo [37, 38, 39] and Climax Foods successfully deploy AI to mine ingredient databases, combine plant ingredients, and design formulations that replicate animal textures.
@@
AI accelerates prototyping by predicting textural outcomes before physical trials and reduces time- and resource-intense trial-and-error cycles. In addition, ingredient design and process optimization can balance target textures with environmental impact, for example, to reduce greenhouse gas emissions. Another important application is personalization, for example designing softer foods for elderly populations to improve swallowing safety. LLMs can also convert consumer feedback such as too chewy or not crispy enough [45] into quantifiable parameters for product adjustment, and bridge the gap between perception and engineering [69]. To address data sparsity, LLMs can generate synthetic training data by pairing realistic texture descriptors with mechanical properties. Ideally, AI can integrate data across multiple length scales, from molecular interactions to macroscopic structure, to capture the hierarchical origins of texture. Inversely, future machine learning tools could help reverse engineer ingredients and process conditions to achieve a desired texture. More broadly, generative design tools, similar to those used for inorganic materials [7], hold the potential to propose ingredient–process combinations that meet desired textural constraints.
@@
Significant obstacles remain before AI-driven texture engineering will truly become routine. The biggest challenge is the lack of large, open datasets that couple textural, sensory, and processing data. Industrial data sharing is rare and data fusion remains technically challenging, as mechanical, rheological, visual, and compositional data are inherently multimodal and often unstructured. The ultimate vision would be an AI system for inverse design–with a specific target texture as input and optimal ingredients and processes as output–that transforms texture development into a predictive, model-guided science.
@@
Sensory: Learning perceptual representations. Ultimately, food is not a material design challenge, but a sensory experience. Aroma, flavor, and texture are the strongest determinants of consumer acceptance. The central challenge is to map measurable physical and chemical properties of ingredients to the multidimensional space of human perception. Machine learning provides a framework to learn this mapping and enables both the prediction of sensory profiles and the computational design of novel ingredients with desired characteristics (Fig. 3D). But how far are current AI models from accurately predicting human sensory perception and from generating novel flavor and aroma experiences?
@@
Traditional sensory science has relied on human panels for perceptual evaluation and on analytical instruments such as gas chromatography–mass spectrometry for chemical composition. Panels remain the gold standard, but are slow and expensive, while instruments offer precision without perceptual insight. Early machine learning approaches combined biomimetic sensors, such as E-noses, with classifiers [70], but required heavy feature engineering and were limited to narrow classification tasks [71]. The field has since shifted toward deep representation learning [72], which learns general-purpose sensory maps directly from data. The Principal Odor Map exemplifies this trend: a graph neural network organizes molecules into a vector space where distance encodes perceptual similarity with human-level accuracy in odor labeling: On a prospective validation set of 323 novel odorants, the model’s predictions closely matched a trained panel mean with per-molecule correlations approaching R $\approx$ 0.6 in representative cases [22]. Researchers now employ attention-based architectures to extend this mapping to molecular mixtures [73]. Algorithmic tools for predictive sensory modeling are maturing rapidly, but progress is constrained by the scarcity and cost of high-quality sensory data.
@@
Learned sensory representations open powerful opportunities for food innovation. High-throughput virtual screening of molecular libraries can identify novel flavor compounds for plant-based meats and accelerate product development [11]. Integrating predictive models with low-cost, miniaturized sensors enables real-time quality monitoring in manufacturing and packaging [74], for example with wireless freshness indicators that reduce food waste [75]. These tools also provide systematic frameworks for biomimicry to predict how modifications of ingredients or processes alter the final sensory profile. Still, a feedback loop remains essential: computational predictions must be continuously validated and refined with real-world experiments to ensure consumer relevance.
@@
Today, the bottleneck is no longer algorithmic capacity but high-quality data acquisition [76]. Deep learning models require large cross-scale datasets, yet sensory panels cannot generate data at scale to capture the vast chemical space of food [45]. The path forward lies in closing the loop between modeling and experiment. Generative models such as transformers and diffusion architectures are algorithmically ready, but their success depends on new data pipelines. Developing robust, low-cost hardware and embedding it into automated platforms will be critical to realize the full potential of AI for sensory science. The ultimate vision is a self-driving sensory lab that integrates chemical synthesis, sensory characterization, and machine learning to autonomously generate, predict, and validate flavor, aroma, and texture perception across diverse food systems.
@@
Manufacturing: Bridging lab and industrial scale. Even the best foods will fail if they cannot be produced consistently and affordably. Manufacturing convincing meat analogues is inherently complex because technical goals such as texture, flavor, appearance, and shelf life interact non-linearly with environmental and economic constraints. Relevant datasets span composition, structure, rheology, thermal behavior, processing steps such as extrusion or fermentation, as well as telemetry, sensory, and consumer data. Through digital twins, computer vision, and other modes, AI is emerging as a key technology to integrate these multimodal datasets and streamline process optimization. Typical model inputs include formulation and process profiles, while outputs can range from texture, color, and flavor to cost and environmental impact. But can these models deliver reliable, scalable, and cost-effective production paths without compromising sensory quality or sustainability goals?
AI and ML span across the manufacturing pipeline, from formulation and process design to in-line quality control. High-moisture extrusion produces fibrous structures that mimic muscle, but outcomes depend on moisture, temperature, and screw configuration. Bayesian optimization explores this space more efficiently than trial-and-error methods [78]. In a pilot study, it identified optimal parameters in ten trials versus fifteen for response-surface methodology and reduced validation errors from 61% to below 15% [79]. Digital twins support soft sensing and control, but most applications remain at pilot scale [81]. Data fragmentation persists across proprietary logs and repositories with inconsistent metadata. To address these limitations, FAIR data principles [82] and LLM-based extraction [83] have successfully employed knowledge graphs to link formulation, process, structure, and sensory performance.
@@
AI is opening multiple opportunities for advancing alternative protein manufacturing. Engineers can use machine learning and Bayesian optimization to tune high-moisture extrusion parameters to target fibrousness, yield, and energy efficiency, while digital twins evaluate what-if simulations for scale-up and real-time sensing [77, 78, 80]. Food scientists can use informatics pipelines to rank plant proteins and design blends at scale using rheology–functionality maps that allow ingredient discovery well beyond commodity sources [79, 82]. LLMs further act as formulation copilots, transform unstructured texts into structured knowledge graphs, and suggest experiments to resolve uncertainty [82]. At the same time, computer vision and hyperspectral analytics can advance in-line quality assurance by linking color and structure to sensory proxies to enable continuous assessment during production [80]. Together, these approaches accelerate prototyping, reduce waste, and bridge lab and industrial manufacturing.
@@
Scaling these tools requires high-quality, shareable datasets that link formulations, processes, structures, and sensory outcomes with standardized metadata. Proprietary and fragmented data currently hinder model transfer and benchmarking and underscore the need for domain-specific FAIR implementations, shared testbeds, and community baselines [81]. Ingredient variability remains a crucial bottleneck and robust hybrid models are needed to achieve reliable process control [77, 78]. Industrial adoption also demands capital investment, workforce training, and validation on real production lines since most food digital twins have not been tested at scale [80]. Shared data standards and LLM-enabled knowledge bases can bridge lab and industrial scale and enable consistent, cost-efficient manufacturing at or below price parity.
@@
Recipes: Personalizing food experiences. Beyond production, food preparation shapes consumer experience. Individuals and foodservice operations must balance enjoyment with health, cost, convenience, and allergens while advancing broader goals such as human and planetary health. LLMs offer strong potential because they draw on large recipe datasets and integrate sources such as Recipes1M, USDA FoodData Central, the HESTIA climate database, and Food.com reviews. The key challenge is to embed human knowledge so that models generalize across diverse users, dietary needs, and cultural traditions.
Recent research evaluates and refines LLMs for recipe generation, revision, and preference prediction. Progress relies on fine-tuning, retrieval-augmented generation, and integration with optimization tools. Benchmarks on Recipe1M show that off-the-shelf models produce unreliable outputs [84]. Fine-tuning improves performance [85] and supports revision tasks [86]. A generative model trained on 2,216 curated burger recipes explored $8.9\times 10^{43}$ candidate formulations and identified Pareto-efficient recipes that reduced environmental impact without compromising taste [44]. In blinded restaurant surveys, some AI-generated burgers outperformed the BigMac benchmark in flavor, texture, and overall liking [87]. Other work demonstrates retrieval-augmented generation for recipes [88] and sustainable menu design [89]. Some LLMs now match expert food scientists in preference modeling [89]. Combined with optimization algorithms, these models can generate recipes that meet nutritional and sustainability targets.
@@
The ability to generate recipes under explicit constraints creates opportunities across individual and institutional contexts. For individuals, fine-tuned LLMs could power AI health coaches that teach users how to modify meals to be more sustainable, nutritious, or culturally appropriate, while respecting restrictions such as allergies or religious requirements. Chefs and students could practice recipe generation in educational applications, receiving AI-driven feedback to refine their skills [90]. In institutional settings, LLMs could support foodservice operations in offices, universities, or hospitals, tailoring descriptions, ingredients, and preparation methods under strict constraints. This is particularly valuable in programs such as “Food as Medicine,” which deliver cost- and time-sensitive meals for patients with chronic conditions after hospital discharge [91]. In all these scenarios, AI could accelerate constrained recipe generation, broaden access to healthy and sustainable foods, and scale public health interventions that rely on tailored meal planning.
@@
Despite this promise, important challenges remain. Off-the-shelf LLMs perform unevenly across cultures, fail to capture diverse preferences, and produce numerically inconsistent or unsafe outputs. They lack molecular-level understanding and have limited knowledge of food safety. Training models from scratch requires substantial resources. These limitations introduce risks of impractical or unsafe recipes and increase environmental cost. Mitigation strategies include integrating external datasets, using open-source pre-trained models, and coupling LLMs with optimization tools. Beyond technical concerns, systems must preserve user agency and support human expertise. A more human-centered approach would allow LLMs to assist with drafting and revision, while users adapt recipes to their needs, preferences, and cultures.
Conclusions